OpenStudy (anonymous):

Help please Solve the given equation. sin2 θ − 5 sin θ − 6 = 0

OpenStudy (doc.brown):

Do you mean\[2\sin\theta-5\sin\theta-6=0\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If that's the case, then you can just add the sine terms and solve for theta \[2 \sin \theta - 5 \sin \theta - 6=0\] \[-3 \sin \theta - 6 =0\] If the expression is \[ \sin ^2 \theta - 5\sin \theta - 6 = 0\] you can treat "sin theta" like an x in the normal polynomial \[x^2-5x-6=0\] and use the quadratic formula to find two values for sin theta \[ \sin\theta = \frac{-b±\sqrt{b^2-4ac}}{2a} = \frac{-(-5)±\sqrt{(-5)^2-4(1)(-6)}}{2(1)}\]